# sylinline,color=cyan

### 23 Mar 2026 v1

@@ 4.2. The smooth critical locus
Now we prove Theorem 4.1.
Proof of Theorem 4.1. By Proposition 4.7, we have
@@
In the case where the projection is non-dominant, Theorem 4.1 recovers the result in [TEV05, Theorem 1.21].
@@ 5. Purity of smooth discriminant
In this section, we will prove Theorem 1.4. Again, we assume $X\subseteq\mathbb{P}^{N}$ is an irreducible projective variety over an algebraically closed field with char 0. We work with the notion of smooth discriminant (cf. Definition 3.1) for generic projection of $X$. When the projection is finite and dominant, we prove Theorem 5.7, where the discriminant is clean.
@@ 5.1. Factorization of projections
First, we show a direct consequence of Theorem 4.1, which characterize the discriminant locus $D$ inductively as a discriminant locus of projection to a general hyperplane:
@@ 5.2. Proof of purity theorem
If $\textup{codim}(X^{\perp})>k$, the $X^{\perp}$ does not intersect with a general linear space of dimension $k$, hence by Theorem 1.3, the discriminant locus is empty; If $\textup{codim}(X^{\perp})=k$, the intersection $X^{\perp}$ with a general $\mathbb{P}^{k}$ consists of finitely many points, whose dual variety is union of hyperplanes; If $\textup{codim}(X^{\perp})<k$, then Bertini theorem implies $\mathbb{P}^{k}\cap X^{\perp}$ is irreducible, hence by Theorem 1.3, its dual variety, the smooth discriminant $\Delta$ is irreducible as well.
@@ 5.3. Purity of smooth critical locus
With the same assumption as before, then the map $\pi|_{(X|_{V})^{\perp}}$ is generically one-to-one. In particular, for generic projection, the map from the smooth critical locus to the smooth discriminant locus
@@ 5.4. Finite branched cover
Now we specialize Theorem 4.1 to the case where the generic projection is finite and dominant.
@@
be a generic projection. Note that this is regular and is a finite branched cover. Then the branch divisor $\Delta=\Delta_{Fitt}$ is the discriminant of $\pi$. Moreover,
@@ 6.1. Empty smooth discriminant
When $X$ is smooth, then $\tilde{\pi}$ is a smooth proper morphism iff $X$ is $\textup{codim}(X^{\perp})>\dim(\mathbb{P}V)$.
@@ 7. Examples and Applications
In this section, we illustrate the applications of the main theorems 1.3 and 1.4 in three examples of surfaces and threefold. In Section 7.1, we take successive generic projections of Veronese surface to hyperplanes. Correspondingly in the dual space, we find special semistable cubic hypersurfaces that are related by taking successive general hyperplane sections; In Section 7.2, we compute the number of nodes and cusps of the discriminant curve arising from the generic projection of a smooth projective surface and of the Veronese threefold.
@@ 7.2. Discriminant curves and their singularities
We may assume $X$ is not dual defective, then the discriminant locus $\Delta$ is an irreducible curve (cf. Theorem 1.4). The singularities $\Delta$ are expected to only have ordinary nodes and cusps. Under this assumption, the numbers of nodes $\delta$ and cusps $\kappa$ can be explicitly deduced from the duality theorem 1.3 and Plücker formulas.
@@ 7.2.1. Generic projection of surfaces
We revisit the discriminant invariants of generic projection of a surface computed in [CMT01], which are used to describe the a special type III degeneration of K3 surfaces. Here we give a different proof of the results using projective duality.
@@
The branch cover $S\to\mathbb{P}^{2}$ has ramification locus $\tilde{\Delta}$. In [CF11], it shows that the map $\tilde{\Delta}\to\Delta$ is a desingularization, and the discriminant curve $\Delta$ has only nodes and cusps as its singularities. Here we assume this fact.
@@ 7.2.2. A net of quadrics and an elliptic threefold
By Theorem 1.3, the discriminant curve $\Delta$ of $X\dashrightarrow\mathbb{P}^{2}$ is the dual curve of a general plane section of the dual variety $\Delta^{\perp}=(\mathbb{P}^{2})^{\vee}\cap X^{\perp}$. Suppose $\Delta^{\perp}$ has only nodes and cusps, then their numbers $\delta^{\perp}$ and $\kappa^{\perp}$ are computed in [DL81, (2.2), (2.3)], [ROB76]. Then $g(\Delta^{\perp})=g(\Delta)$, $\deg(\Delta^{\perp})$, $\delta^{\perp}$ and $\kappa^{\perp}$ are related to $\deg(\Delta)$, $\delta$, $\kappa$ via the Plücker formulas.
@@ 8.3. Classical results
9. Monodromy and Braid group
@@
Let $\textup{Sym}^{k}(\mathbb{P}^{1})$ be the symmetric product of $\mathbb{P}^{1}$. Let $D^{k}=\{[x_{1},\dotsb,x_{k}]\in\textup{Sym}^{k}(\mathbb{P}^{1})|\exists\ 1\leq i<j\leq k,\ x_{i}=x_{j}\}$ be the fat diagonal. Then the braid group $B_{k}$ of $k$ points on $\mathbb{P}^{1}$ is the fundamental group
@@ 9.1. Braid monodromy from generic finite projection
Write $C=Z\cap p^{-1}(\ast)$. Note that the restriction $p^{\circ}\mathrel{\mathop{\ordinarycolon}}Z\setminus p^{-1}(\Delta_{Fitt})\to\mathbb{P}^{1}\setminus\Delta_{Fitt}$ is étale, so the fiber over $t$ is an unordered configuration of $m$ distinct points in the line $\mathbb{P}^{1}$. This induces a continuous map $\mathbb{P}^{1}\setminus\Delta_{Fitt}\to\textup{Sym}^{m}(\mathbb{P}^{1})\setminus D^{m}$, hence a group homomorphism
@@
When $Z$ is smooth and the projection is general, the projection $Z\to\mathbb{P}^{1}$ is simply branched, and $\Delta_{Fitt}=\Delta$ consists of branch points. The monodromy action of each generating loop of $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)$ is given by Dehn twist along a pair of points. As the covering space $p^{\circ}$ is connected, the monodromy action on $C$ is transitive, and there exists an ordering $C=\{z_{0},\dotsb,z_{m-1}\}$ and generating loops $\gamma_{0},\dotsb,\gamma_{m-1}$ of $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)$ such that $\gamma_{i}$ is the Dehn twist along the pair $(z_{i\mod(m-1)},z_{i+1\mod(m-1)})$. It is classically known that the action of the subgroup $\langle\gamma_{0},\dotsb,\gamma_{m-1}\rangle$ on $p^{-1}(\ast)\setminus Z$ is exactly the braid relation [ART25], and therefore the image of $\langle\gamma_{0},\dotsb,\gamma_{m-1}\rangle$ under the map $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)\to B_{m}$ coincides with the braid group, so the map is surjective. ∎
@@
Note that there is a surjection $B_{m}\to\operatorname{\mathrm{Mod}}(\mathbb{P}^{1}\setminus C)$ to the mapping class group of $\mathbb{P}^{1}\setminus C$ (see [FM12, Section 9.1]). The composition of $g$ with this map is the induced monodromy action on $\mathbb{P}^{1}\setminus C$.
@@
    If $Y$ is normal, then $\varphi$ is surjective, and $\Delta_{Fitt}=\Delta$ is an irreducible hypersurface of degree m(m1)m(m-1).
@@
For part (2), Suppose $Y$ is normal, Then by Lemma 3.6 implies that $\Delta_{Fitt}=\Delta$ coincides with the smooth discriminant. For a general line $\mathbb{L}=\mathbb{P}^{1}\subseteq\mathbb{P}^{N}$, the preimage with $p^{-1}(\mathbb{L})\mathrel{\mathop{\ordinarycolon}}=Z$ is a general plane section of $Y$. Since $Y$ is normal and smooth in codimension one, $Z$ is a smooth plane curve of degree $m$. The projection $\mathbb{P}^{2}\dashrightarrow\mathbb{L}$ restricts to a finite branched cover
@@
and hence surjectivity of $\pi_{1}\left(\mathbb{L}\setminus\Delta,\ast\right)\twoheadrightarrow B_{m}$ implies the surjectivity of the second map.
For the last part, we note $\deg(\Delta)$ is the number of the (simple) branch points of $Z\to\mathbb{P}^{1}$. Specifically, it follows from Theorem 8.7 that
@@
is the monodromy representation permutating the $m$ points.
In the set-up of In Proposition 9.3 (1) and without additional assumption, one can still show $\rho$ is surjective. This is an application of the Uniform Principle Theorem of an irreducible curve [ACG+85, p.111].
9.2. Varieties with normal duals
Proposition 9.5.
$\pi_{1}(\mathbb{P}^{N}\setminus X^{\perp},*)\twoheadrightarrow B_{m}.$
Proof.
By Theorem 5.7, the dual variety of $\mathcal{X}\cap H$ is the discriminant of generic projection $\mathcal{X}^{\perp}\to(\mathbb{P}^{N})^{\vee}$. The rest follows from Proposition 9.3. ∎
@@ 9.2. Varieties with normal duals
The first part follows from Proposition 9.3 and the fact that $\Delta=\Delta_{Fitt}$ is the branch locus of the finite cover $Y\to\mathbb{P}^{N}$. The second part follows from Proposition 9.5. ∎
@@
Let $\mathcal{X}=v_{2}(\mathbb{P}^{n})$ be the 2nd Veronese embedding of $\mathbb{P}^{n}$. Then $\mathcal{X}^{\perp}$ is the space of symmetric $(n+1)\times(n+1)$ matrices with determinant zero, and therefore has degree $n+1$. The hyperplane section $X=\mathcal{X}\cap H$ is isomorphic to a quadric hypersurface in $\mathbb{P}^{n}$ embedded in $\mathbb{P}^{N}$ via $v_{2}$. The dual variety $X^{\perp}$ is a hypersurface of degree $n(n+1)$. The braid monodromy induces surjection
@@
Remark 9.7.
Note that by Zariski’s theorem [VOI07, Theorem 3.22], the fundamental group of the complement of the hypersurface $X^{\perp}$ is isomorphic to $\pi_{1}((\mathbb{P}^{2})^{\vee}\setminus X^{\perp}\cap(\mathbb{P}^{2})^{\vee},*)$, the fundamental group of the complement of a plane section of $X^{\perp}$. According to Theorem 4.1, the plane curve $X^{\perp}\cap(\mathbb{P}^{2})^{\vee}$ is isomorphic to the smooth discriminant of the generic projection $X\dashrightarrow\mathbb{P}^{2}$. Suppose that it only has nodes and cusps, one can deduce their number using the techniques in Section 7.2.

### 19 Apr 2026 v2

@@ 4.2. The smooth critical locus
4.3. Proof of Theorem 4.1.
By Proposition 4.7, we have
@@
Remark 4.9.
An analogous statement of Proposition 4.7 was obtained in [FT18, Proposition 3.9] in the generic linear projection of the affine-analytic setting. In our projective setting (and as well Theorem 4.1), the biduality isomorphism $N_{(\mathbb{P}^{N})^{\perp}}^{\perp}(X^{\perp})\cong N_{\mathbb{P}^{N}}^{\perp}X$ is necessary in the proof.
4.4. Relation to other work
In the case where the generic projection is non-dominant, Theorem 4.1 recovers the result in [TEV05, Theorem 1.21].
@@ 4.4. Relation to other work
When $X$ is a hypersurface, Theorem 4.1 recovers a result of Piene [PIE78, Corollary 4.3].
@@ 5. Purity of smooth discriminant
In this section, we will prove Theorem 5.3, which implies Theorem 1.4. Again, we assume $X\subseteq\mathbb{P}^{N}$ is an irreducible projective variety over an algebraically closed field with char 0. We work with the notion of smooth discriminant (cf. Definition 3.1) for generic projection of $X$. When the projection is finite and dominant, we prove Theorem 5.7, where the discriminant is clean.
@@ 5.1. Factorization of projections
First, we show a direct consequence of Theorem 4.1, which characterize the discriminant locus $D$ inductively as a discriminant locus of projection to a general hyperplane. This phenomenon was also observed in [TT88, Proposition 2.2.5]. One refers to 7.1 for an explicit example for this point of view.
@@ 5.2. Proof of purity theorem
If $\textup{codim}(X^{\perp})>k$, $X^{\perp}$ does not intersect with a general linear space of dimension $k$, hence by Theorem 1.3, the discriminant locus is empty; If $\textup{codim}(X^{\perp})=k$, the intersection $X^{\perp}$ with a general $\mathbb{P}^{k}$ consists of finitely many points, whose dual variety is a union of hyperplanes; If $\textup{codim}(X^{\perp})<k$, then Bertini theorem implies $\mathbb{P}^{k}\cap X^{\perp}$ is irreducible, hence by Theorem 1.3, its dual variety, the smooth discriminant $\Delta$ is irreducible as well.
@@ 5.3. Purity of smooth critical locus
With the same assumption as before, the map $\pi|_{(X|_{V})^{\perp}}$ is generically one-to-one. In particular, for generic projection, the map from the smooth critical locus to the smooth discriminant locus
@@ 5.4. Finite branched cover
Now we specialize Theorem 4.1 to the case where the generic projection is finite and dominant. Recall that in this case, the smooth discriminant coincides with Fitting discriminant (See Lemma 3.6).
@@
be a generic projection. Note that this is regular and is a finite dominant morphism. Then the branch divisor $\Delta=\Delta_{Fitt}$ is the discriminant of $\pi$. Moreover,
@@ 6.1. Empty smooth discriminant
When $X$ is smooth, then $\tilde{\pi}$ is a smooth proper morphism if and only if $\textup{codim}(X^{\perp})>\dim(\mathbb{P}V)$.
@@ 7. Examples and Applications
In this section, we illustrate the applications of the main theorems 1.3 and 1.4 in three examples involving surfaces and threefolds. In Section 7.1, we take successive generic projections of Veronese surface to hyperplanes. Correspondingly in the dual space, we find special semistable cubic hypersurfaces that are related by taking successive general hyperplane sections; In Section 7.2, we compute the number of nodes and cusps of the discriminant curve arising from the generic projection of a smooth projective surface and of the Veronese threefold.
@@ 7.2. Discriminant curves and their singularities
When $X$ is not dual defective, then the discriminant locus $\Delta$ is an irreducible curve (cf. Theorem 1.4). The singularities $\Delta$ are expected to only have ordinary nodes and cusps. Under this assumption, the numbers of nodes $\delta$ and cusps $\kappa$ can be explicitly deduced from the duality theorem 1.3 and Plücker formulas.
@@ 7.2.1. Generic projection of surfaces
We revisit the discriminant invariants of generic projection of a surface computed in [CMT01], which are used to describe a special type III degeneration of K3 surfaces. Here we give a different proof of the results using projective duality.
@@
The branch cover $S\to\mathbb{P}^{2}$ has ramification locus $\tilde{\Delta}$. In [CF11], it shows that the map $\tilde{\Delta}\to\Delta$ is a desingularization, and the discriminant curve $\Delta$ has only nodes and cusps as its singularities. Here we take this fact.
@@ 7.2.2. A net of quadrics and an elliptic threefold
By Theorem 1.3, the discriminant curve $\Delta$ of $X\dashrightarrow\mathbb{P}^{2}$ is the dual curve of a general plane section of the dual variety $\Delta^{\perp}=(\mathbb{P}^{2})^{\vee}\cap X^{\perp}$. Suppose $\Delta^{\perp}$ has only nodes and cusps, then their numbers $\delta^{\perp}$ and $\kappa^{\perp}$ are computed in [DL81, (2.2), (2.3)], [ROB76]. Then $g(\Delta^{\perp})=g(\Delta)$, $\deg(\Delta^{\perp})$, $\delta^{\perp}$ and $\kappa^{\perp}$ are related to $\deg(\Delta)$, $\delta$, $\kappa$ via the Plücker formulas. In general, this method can characterize codimension-one singularities of discriminant $\Delta$ for generic projection $X\dashrightarrow\mathbb{P}^{k}$ for any $k$.
@@ 8.3. Classical results
9. Monodromy and braid groups
@@
Let $\textup{Sym}^{k}(\mathbb{P}^{1})$ be the symmetric product of $\mathbb{P}^{1}$. Let $D^{k}=\{[x_{1},\dotsb,x_{k}]\in\textup{Sym}^{k}(\mathbb{P}^{1})|\exists\ 1\leq i<j\leq k,\ x_{i}=x_{j}\}$ be the fat diagonal. Then the (spherical) braid group $B_{k}$ of $k$ points on $\mathbb{P}^{1}$ is the fundamental group
@@ 9.1. Braid monodromy from generic finite projection
Write $F=p^{-1}(\ast)$. Note that the restriction $p^{\circ}\mathrel{\mathop{\ordinarycolon}}Z\setminus p^{-1}(\Delta_{Fitt})\to\mathbb{P}^{1}\setminus\Delta_{Fitt}$ is étale, so the fiber over $t$ is an unordered configuration of $m$ distinct points in the line $\mathbb{P}^{1}$. This induces a continuous map $\mathbb{P}^{1}\setminus\Delta_{Fitt}\to\textup{Sym}^{m}(\mathbb{P}^{1})\setminus D^{m}$, hence a group homomorphism
@@
When $Z$ is smooth and the projection is general, the projection $Z\to\mathbb{P}^{1}$ is simply branched, and $\Delta_{Fitt}=\Delta$ consists of branch points. The monodromy action of each generating loop of $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)$ is given by Dehn twist along a pair of points. As the covering space $p^{\circ}$ is connected, the monodromy action on $F$ is transitive, and there exists an ordering $F=\{z_{0},\dotsb,z_{m-1}\}$ and generating loops $\gamma_{0},\dotsb,\gamma_{m-1}$ of $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)$ such that $\gamma_{i}$ is the Dehn twist along the pair $(z_{i\mod(m-1)},z_{i+1\mod(m-1)})$. It is classically known that the action of the subgroup $\langle\gamma_{0},\dotsb,\gamma_{m-1}\rangle$ on $p^{-1}(\ast)\setminus Z$ is exactly the braid relation [ART25], and therefore the image of $\langle\gamma_{0},\dotsb,\gamma_{m-1}\rangle$ under the map $\pi_{1}\left(\mathbb{P}^{1}\setminus\Delta,\ast\right)\to B_{m}$ coincides with the braid group, so the map is surjective. ∎
@@
Note that there is a surjection $B_{m}\to\operatorname{\mathrm{Mod}}(\mathbb{P}^{1}\setminus F)$ to the mapping class group of $\mathbb{P}^{1}\setminus F$ (see [FM12, Section 9.1]). The composition of $g$ with this map is the induced monodromy action on $\mathbb{P}^{1}\setminus F$.
@@
    Assume moreover that $Y$ is normal. Then $\Delta_{Fitt}=\Delta$, and
    • (a)
      the induced braid monodromy representation
      $\varphi\mathrel{\mathop{\ordinarycolon}}\pi_{1}(\mathbb{P}^{N}\setminus\Delta,\ast)\to B_{m}$
      is surjective;
    • (b)
      $\Delta$ is a hypersurface of degree
      $\deg(\Delta)=m(m-1);$
    • (c)
      if $\mathcal{X}=Y^{\perp}\subseteq(\mathbb{P}^{N+1})^{\vee}$ and $X=H\cap\mathcal{X}$ is the general hyperplane section determined by the projection, then
      $\Delta\cong X^{\perp};$
    • (d)
      if $N>1$, then $\Delta$ is irreducible.
@@
For part (2), Suppose $Y$ is normal, Then Lemma 3.6 implies that $\Delta_{Fitt}=\Delta$ coincides with the smooth discriminant. For a general line $\mathbb{L}\subseteq\mathbb{P}^{N}$, let $Z\mathrel{\mathop{\ordinarycolon}}=p^{-1}(\mathbb{L})$. Then $Z$ is a general plane section of $Y$. Since $Y$ is normal and smooth in codimension one, Bertini theorem implies that $Z$ is a smooth plane curve of degree $m$. The projection $\mathbb{P}^{2}\dashrightarrow\mathbb{L}$ restricts to a finite branched cover
@@
and hence surjectivity of $\pi_{1}\left(\mathbb{L}\setminus\Delta,\ast\right)\twoheadrightarrow B_{m}$ implies the surjectivity of the second map. This proves (a)
For (b), we note $\deg(\Delta)$ is the number of the (simple) branch points of $Z\to\mathbb{P}^{1}$. Specifically, it follows from Theorem 8.7 that
@@
The statement (c) follows from Theorem 5.7. For irreducibility of $\Delta$, by Theorem 5.7 it suffices to show that $\dim Y^{\perp}>1$. Suppose instead that $Y^{\perp}$ is a curve. Let $S$ be a general $\mathbb{P}^{3}$-section of $Y$. Since $Y$ is normal, $S$ is a normal surface. By Theorem 4.1, the surface $S$ is the dual variety of the image of a general projection of $Y^{\perp}$ to $(\mathbb{P}^{3})^{\vee}$. But the dual variety of a nondegenerate space curve is a tangential surface, hence is singular along the osculating curve ([DOL12, Proposition 10.24]), contradicting the normality of $S$. ∎
Remark 9.4.
Note that by Zariski’s theorem on fundamental group of complement of hypersurface [VOI07, Theorem 3.22], the $\pi_{1}$ above is isomorphic to fundamental group of complement of plane curve $\pi_{1}(\mathbb{P}^{2}\setminus C,*)$, where $C=X^{\perp}\cap\mathbb{P}^{2}$ is a general plane section of degree m(m1)m(m-1). In particular, when $m\geq 4$, Proposition 9.3 (2a) implies that $\pi_{1}(\mathbb{P}^{2}\setminus C,*)$ contains a free group with two generators, and is "big" in the sense of [DOZ98]. Note that the plane curves $C$ here are in general different from the examples studied in [DOZ98].
@@
is the monodromy representation permuting the $m$ points.
In the set-up of Proposition 9.3 (1) and without additional assumption, one can still show $\rho$ is surjective. This is an application of the Uniform Principle Theorem of an irreducible curve [ACG+85, p.111].
@@
It follows from Proposition 9.3 (2). ∎
9.2. Varieties with normal duals
The following is a direct consequence of Proposition 9.3 by assuming $\mathcal{X}$ is smooth. This proves Corollary 1.9.
Proposition 9.6 (Corollary 1.9).
$\pi_{1}((\mathbb{P}^{N})^{\vee}\setminus X^{\perp},*)\twoheadrightarrow B_{m}.$
@@ 9.2. Varieties with normal duals
Let $\mathcal{X}=v_{2}(\mathbb{P}^{n})$ be the second Veronese embedding of $\mathbb{P}^{n}$. Then $\mathcal{X}^{\perp}$ is the determinantal hypersurface of singular symmetric $(n+1)\times(n+1)$ matrices, so $\deg(\mathcal{X}^{\perp})=n+1$. For a general hyperplane section $X=\mathcal{X}\cap H$, the dual variety $X^{\perp}$ is a hypersurface of degree $n(n+1)$, and the braid monodromy induces a surjection
@@
Example 9.8.
In the special case $n=2$, $\mathcal{X}=v_{2}(\mathbb{P}^{2})$ and $\mathcal{X}^{\perp}$ is a normal cubic hypersurface. (See Section 7.1.) A general hyperplane section $X=\mathcal{X}\cap H$ is the rational normal quartic, and $X^{\perp}$ is the discriminant hypersurface in $\textup{Sym}^{4}(\mathbb{P}^{1})\cong\mathbb{P}^{4}$. A general plane section of $X^{\perp}$ is the classical rational sextic with $4$ nodes and $6$ cusps studied by Zariski [ZAR36]. Its complement has fundamental group $B_{4}$. Hence in this case the braid monodromy surjection $\pi_{1}((\mathbb{P}^{4})^{\vee}\setminus X^{\perp},*)\twoheadrightarrow B_{3}$ is not an isomorphism.